OpenStudy (anonymous):

Suppose 5 men and 7 women are on a crowded elevator. At the next floor, four people get off the elevator. Find the probability that three are women.

OpenStudy (anonymous):

the number of ways to choose four people out of a set of 12 is \[\binom{12}{4}\] that is your denominator

OpenStudy (anonymous):

3 are women, one is a man, \[\binom{7}{3}\times \binom{5}{1}\] that is your numerator

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know how to compute \[\binom{7}{3}\]? sometimes this is written as \(_7C_3\)

OpenStudy (anonymous):

@misty1212 yes, i know it as 7C3 would it be \[7c3\div12c4\]

OpenStudy (misty1212):

no, you need another 5 up top, because there are five men and one gets off too

OpenStudy (misty1212):

\[\frac{_7C_3\times 5}{_{12}C_4}\]

OpenStudy (anonymous):

@misty1212 thank you so much c: that helps a lot

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]